the length of a rectangle is 3 meters less than tiwice the width.if the area is 464 square meters,find the dimensions 464
Answer by rfadrogane(214)

the length of a rectangle is 3 meters less than twice the width.if the area is 464 square meters,find the dimensions 464.
Sol'n:
Let l-the length
w-the width
l=2w-3 ----(1)
lw=464 ----(2)
substitute (1) into (2)
w(2w-3)=464
2w^2-3w-464=0
factor:
(w-16)(2w+29)=0
take w = 16m, then substitute to (1).
l=2(16)-3 = 29
thus, the dimension is
29m x 16m -----answer
^_^